OpenStudy (anonymous):

Calculus Easy Problems!!Those are easy problems but I don't know why to get start with the the numbers!! 1.If f(2)=14, f′ is continuous, and ∫(upper limit:7,lower limit:2)f′(t)dt=25, what is the value of f(7)? 2.If g(1)=−3, g(5)=9, and ∫(upper limit:5,lower limit:1)g(x)dx=−5, evaluate the integral ∫(upper limit:5,lower limit:1)xg′(x)dx.

OpenStudy (anonymous):

it IS easy, you just don't know how to do it. Everything is easy WHEN you know how. Stop stressing. the answer to the definite integral is f(upper limit) - f(lower limit) = number so f(7) - f(2) = 25, and you're given f(2) = 14. Find f(7). there now... that's easy isn't it?

OpenStudy (anonymous):

I got the first one!!thank you!!

OpenStudy (anonymous):

you're welcome :)

OpenStudy (dumbcow):

For second one use integration by parts u =x dv = g'(x) du =dx v = g(x) =x*g(x) -integral g(x) integral g(x) = -5 = x*g(x) +5 evaluate x*g(x) at 1 and 5

OpenStudy (anonymous):

loads of thanks!!